# ToMiE: Towards Modular Growth in Enhanced SMPL for 3D Human Gaussians with Animatable Garments

### v1

@@ Abstract
In this paper, we highlight a critical yet often overlooked factor in most 3D human tasks, namely modeling humans with complex garments. It is known that the parameterized formulation of SMPL is able to fit human skin; while complex garments, e.g., hand-held objects and loose-fitting garments, are difficult to get modeled within the unified framework, since their movements are usually decoupled with the human body. To enhance the capability of SMPL skeleton in response to this situation, we propose a modular growth strategy that enables the joint tree of the skeleton to expand adaptively. Specifically, our method, called ToMiE, consists of parent joints localization and external joints optimization. For parent joints localization, we employ a gradient-based approach guided by both LBS blending weights and motion kernels. Once the external joints are obtained, we proceed to optimize their transformations in $SE(3)$ across different frames, enabling rendering and explicit animation. ToMiE manages to outperform other methods across various cases with garments, not only in rendering quality but also by offering free animation of grown joints, thereby enhancing the expressive ability of SMPL skeleton for a broader range of applications. The code is available at https://github.com/Yifever20002/ToMiE.
Figure 1: We show two common types of complex garments in everyday life: hand-held objects and loose-fitting garments, which cannot be accurately represented by the standard SMPL model. Our ToMiE can realize adaptive growth to enhance the representation capability of SMPL without needing for time-consuming case-specific customization, achieving state-of-the-art results in both rendering and human (including complex garments) animation.
@@ 1 Introduction
The neural-based 3D human rendering has been empowered by the cutting-edge technique, 3D Gaussian Splatting (3DGS) [18], for its real-time and high-quality novel view synthesis performance. By representing 3D gaussians under the canonical T-pose and utilizing the pre-extracted SMPL skeleton in the observation space, 3D human rendering results can be obtained from novel views in any frame. This stream of approaches proves high quality in rendering 3D humans that conform to the SMPL paradigm (e.g., avatars in tight-fitting clothing). However, we raise concerns regarding its ability to handle human modeling involving daily garments, such as skirts or hand-held objects.
@@
In Fig. 1, we show two cases to illustrate the limitations of current 3D human methods in modeling daily garments. On the one hand, characters shot in-the-wild are dressed in garments with high complexity in dynamics, rather than the tight-fitting garments configured under strict experimental conditions. These loose-fitting garments break the existing methods’ paradigm by assuming that garments should be bounded to the motion warping of SMPL skeleton in the same way as the human body is. On the other hand, the movements of hand-held objects, e.g., mobile phones, are highly decoupled from the human body and thus cannot be represented by the SMPL. Rooted in the aforementioned scenarios, SMPL is observed to suffer from appearance ambiguity when attempting to fit such 3D human models.
@@
To this end, we break through the limitations of modeling 3D human gaussians with complex garments by maintaining an extended joint tree from SMPL skeleton. Although existing SMPL model has the potential capability to manually customize additional skeleton information, time-consuming and case-by-case adjustments are necessary in this case. To overcome this issue, we extend the SMPL skeleton with additional joints for each individual case adaptively. The growth is performed in an explicit and modular manner, and enables the fitting of complex garments, offering high-quality rendering of more complex cases compared to other methods (e.g., novel view synthesis results in Fig. 1).
@@
The main challenge of extending the SMPL skeleton is to determine where and how to grow additional joints. To avoid the unnecessary memory consumption and potential overfitting caused by an arbitrary growth, we first determine which of the joints are supposed to serve as parent joints by a localization strategy. We have empirically observed that parent joints requiring growth wittness larger backpropagation gradients in their associated gaussians due to underfitting. However, determining the association of gaussians with different joints is non-trivial, as the SMPL’s LBS blending weights may not work with the human with garments. To more precisely define such association, we introduce the concept of Motion Kernels based on rigid body priors and combine them with LBS weights, resulting in more accurate gradient-based localization. After growth, we adaptively maintain an extended modular joint tree and update the extra joints by optimizing two MLP decoders for joints positions and rotations. The proposed method, termed as ToMiE, allows for explicit animation of complex garments represented by external joints.
@@
By experiments on complex cases with garments of the DNA-Rendering dataset [5], ToMiE exhibits state-of-the-art rendering quality while maintaining the animatability that is significant for downstream productions. To summarize, our contributions are three-fold as follows:
@@
    ToMiE, a method for creating an enhanced SMPL joint tree via a modular growth strategy, which is able to decouple complex garments from the human body, thereby achieving state-of-the-art results in both rendering and explicit animation on target cases;
@@ 2.1 SMPL-based Human Mesh Avatars
Most of the recent success in digital human modeling can be attributed to the contributions of the SMPL [28] series, which parameterizes the human body as individual shape components and motion-related human poses through 3D mesh scanning and PCA. The pose blend shapes in SMPL describe human body deformations as a linear weighted blending of different joint poses, significantly improving the efficiency of editing and animating digital humans. Furthermore, it has been widely adopted for human body animation, thanks to the methods [6, 39] of estimating SMPL parameters from 2D inputs. Despite their wide range of applications, SMPL and its family still suffer from inherent limitations. Since the originally scanned 3D meshes are skin-tight, according to which the pose blend shapes are learned, the model is unable to handle significantly outlying meshes, such as human with garments like skirts and hand-held objects.
@@ 2.2 Neural Representation for 3D Human
Methods based on neural representations, such as NeRF [30] and 3DGS [18], have also been playing an important part in digital human reconstruction for their high-quality rendering capabilities. Early NeRF-based methods [42, 36, 20, 4, 10, 1, 2, 3, 7, 8, 9] aim to reconstruct human avatars by inputting monocular or multi-view synchronized videos. [40] enforce smooth priors based on neural Signed Distance Function (SDF) to obtain more accurate human geometry. Recent breakthroughs [24, 27, 46, 22, 14, 37, 21, 15, 32, 17, 13, 23, 26, 44, 19, 31, 16, 45] rely on 3DGS, enabling faster and more accurate rendering. All these methods register the T-pose in a canonical space and use LBS weights to guide the rigid transformations. The animation of human body can be achieved by warping points in the canonical space into the observation space that is controlled by per-frame human poses.
2.3 Garments Rendering and Editing in 3D Human
We revisit the methods that consider implicitly improving garment topology. Animatable NeRF [35] defines a per-frame latent code to capture appearance variations across each frame. [21, 11] additionally register global latent bones to compensate for the limitations in garments rendering. They fail to explicitly decouple the garments from human body, making precise control infeasible. Another stream [4, 15] leverages a human poses sequence as contexts to resolve appearance ambiguities. But correspondingly, they require a sequence of human poses for animation, adding to the challenges of editing. Moreover, these methods struggle to fit the object-level pose independent of the human poses sequence, such as hand-held items. We also note that some works [12, 29] introduce diffusion-based generative methods to enhance the realism of garment rendering, but these methods are restricted by the traditional SMPL and overlook the editing of complex garments. Our approach enables explicit decoupling of garments and the human body by extending the SMPL joint tree, allowing for high-quality rendering and explicit animating of garments, including hand-held objects.
@@ 3.1 SMPL(-X) Revisited
The main issue with this LBS-based prior model is that it can only fit tight-fitting avatars conforming to the SMPL(-X) paradigm, making it unsuitable for modeling complex human clothing in more generic cases. It is even more challenging is its inability to handle hand-held objects that are fully decoupled from the human pose. To leverage SMPL(-X) prior without being constrained by garment types, we highlight its extension to complex garment modeling.
@@ 3.2 Human Gaussians Revisited
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (3)
@@ 4 Methods
Figure 2: The pipeline of ToMiE. We initialize the gaussians in the canonical space with a standard SMPL vertices.  (Sec. 4.4) We apply Linear Blend Skinning (LBS) to the gaussian position and utilize a network for rotation and scale correction. During the warmup phase, Adaptive LBS only utilizes the original SMPL skeleton. After adaptive growth, it further includes the newly grown external skeleton. Gaussian rasterization and gradients backpropagation.  (Sec. 4.1, Sec. 4.2) We employ a gradient-based parent joints localization method, and a motion kernel to optimize the gradient assignment process.  (Sec. 4.3) We maintain an extra joint book with MLPs, which generates explicit human pose, enabling the decoupling and explicit animating of complex garments.
Fig. 2 illustrates ToMiE’s modular joint growth and gaussian training strategy. Our goal is to extend the SMPL skeleton to handle complex human garments. However, the abuse of growth can lead to unnecessary computational and memory overheads. For an efficient adaptive growth, we first propose a localization strategy of parent joints to ensure that only necessary joints are grown. Furthermore, we explicitly define the hand-held joints in the $SE(3)$ space and optimize them end-to-end through an MLP, ensuring alignment with the original skeleton of SMPL. The extended skeleton can thus support rendering and explicit animation. To address the limitation of LBS in guiding gaussian attributes apart from the positions, we further fine-tune the rotation and scale during training with a deformation field to achieve better non-rigid warping. Next, we will elaborate on these modules and the training strategy in detail.
@@ 4.1 Motion Kernels-guided Joint Gradient Accumulation
A quantitative metric needs to be identified to determine whether a joint requires growth. We notice that, due to the poor fitting ability of existing SMPL-based human gaussian, it will leave larger gradients in human regions with complex garments (e.g., the hand-held object in Fig. 2 ). In other words, identifying joints with larger accumulated gradients can help to indicate which parent joints are more likely to require growth. Bounded with each gaussian, the accumulated gradients will be first assigned to their corresponding joints. Let $g=\|(g_{x},g_{y},g_{z})\|_{2}$ represent the L2 norm of the gradient at each gaussian position. The gradients accumulation $g_{J}$ for the $k$-th joint can then be computed according to
@@
It is worth emphasizing that $\omega_{k}$ here is a weight term determining the assignment of a gaussian to the $k$-th joint. Under the paradigm of SMPL representation, this weight corresponds to the LBS weight $\omega_{\text{lbs0}}$, and guides the rigid transformation of vertices on the SMPL mesh according to the human pose. To account for the differences between the human mesh with garments and the vanilla SMPL mesh, [14] calculate $\omega_{\text{lbs}}$ by keeping using the LBS weight prior $\omega_{\text{lbs0}}$ and adding an extra learnable network $\Phi_{\text{lbs}}$ for fine-tuning. This formulation (with index $k$ omitted) can be summarized as
@@
where $\text{NN}_{1}$ stands for top-1 nearest-neighbor search algorithm and $\bm{V}$ represents the canonical standard SMPL vertices. It will be, however, clarified by us, that this nearest neighbor-based assigning method will no longer be feasible in cases with complex garments.
@@ 4.4 Inference Process and Training Strategy
In this subsection, we explain how our modular growth is integrated with the training process of human gaussians.
@@
Once the warm-up iterations completes, we begin the modular joint growth. With the MK calculated during the warm-up phase, we can locate the parent joints $\bm{J}_{s}$ that require growth. Then we add grown joints to the extra joint book $B^{e}$, optimizing its positions $\bm{j}^{e}$ and rotations $\bm{r}^{e}$ during the subsequent training. Notably, $\Phi_{\text{lbs}}$ needs to extend its output dimensions to include the blending weights for extra joints, following $K=K^{0}+K^{e}$. Since the extra joints do not have prior LBS weights, their blending weights are entirely learned through $\Phi_{\text{lbs}}$.
@@ 5.1 Dataset
Figure 4: Qualitative comparison on the DNA-Rendering dataset [5]. We show two cases of hand-held objects (0041_10) and loose-fitting garments (0007_04) (from top to bottom). 3DGS-Avatar [37] can fit hand-held objects in some cases (e.g., 0041_10), its pose becomes incorrect due to the lack of joint optimization (cyan lines are parallel). Im4d* [25] achieves high-quality rendering but cannot be animated.
Our method focuses on complex garments and hand-held objects, so we select datasets for experiments accordingly. We notice that the DNA-Rendering dataset [5], by capturing complex scenes of the human body, meets our requirements. Specifically, we selected 8 cases that align with our hypothesis, namely 0041_10, 0090_06, 0176_07, 0800_07 (hand-held objects) and 0007_04, 0014 06, 0051_09, 0811_06 (loose-fitting garments). For each case, we use 24 surrounding views for training and 6 novel surrounding views for testing. All views are synchronized and contained 100 frames each.
In addition to tackle complex garments, it is essential to ensure the model’s performance in typical scenarios involving tight-fitting clothes. Therefore, we additionally test our method on the ZJU-MoCap [36] dataset. Although the tight-clothing cases are too simple to require joint extension, our overall framework still achieves optimal results. Since this part of the experiment is not directly related to modular growth, we refer the readers to the supplemental material for further visualizations.

### v2

# ToMiE: Towards Explicit Exoskeleton for the Reconstruction of Complicated 3D Human Avatars

@@ Abstract
In this paper, we highlight a critical yet often overlooked factor in most 3D human tasks, namely modeling complicated 3D human with with hand-held objects or loose-fitting clothing. It is known that the parameterized formulation of SMPL is able to fit human skin; while hand-held objects and loose-fitting clothing, are difficult to get modeled within the unified framework, since their movements are usually decoupled with the human body. To enhance the capability of SMPL skeleton in response to this situation, we propose a growth strategy that enables the joint tree of the skeleton to expand adaptively. Specifically, our method, called ToMiE, consists of parent joints localization and external joints optimization. For parent joints localization, we employ a gradient-based approach guided by both LBS blending weights and motion kernels. Once the external joints are obtained, we proceed to optimize their transformations in $SE(3)$ across different frames, enabling rendering and explicit animation. ToMiE manages to outperform other methods across various cases with hand-held objects and loose-fitting clothing, not only in rendering quality but also by offering free animation of grown joints, thereby enhancing the expressive ability of SMPL skeleton for a broader range of applications. The code is available at https://github.com/Yifever20002/ToMiE.
Figure 1: We show two complex scenarios in 3D human modeling: with hand-held objects and loose-fitting clothing, which cannot be accurately represented by the standard SMPL model. Our ToMiE can realize adaptive growth to enhance the representation capability of SMPL without needing for time-consuming case-specific customization, achieving state-of-the-art results in both rendering and human (including complex scenarios) animation.
@@ 1 Introduction
The neural-based 3D human rendering has been empowered by the cutting-edge technique, 3D Gaussian Splatting (3DGS) [25], for its real-time and high-quality novel view synthesis performance. By representing 3D gaussians under the canonical T-pose and utilizing the pre-extracted SMPL skeleton in the observation space, 3D human rendering results can be obtained from novel views in any frame. This stream of approaches proves high quality in rendering 3D humans that conform to the SMPL paradigm (e.g., avatars in tight-fitting clothing). However, we raise concerns regarding its ability to handle complicated human modeling involving hand-held objects or loose-fitting clothing.
@@
In Fig. 1, we show two cases to illustrate the limitations of current 3D human modeling. On the one hand, the movements of hand-held objects, e.g., feather duster, are highly decoupled from the human body and thus cannot be represented by the SMPL. On the other hand, characters shot in-the-wild are dressed in clothing with high complexity in dynamics, rather than the tight-fitting clothing configured under strict experimental conditions. These complicated scenarios break the existing methods’ paradigm by assuming that surface of avatars should be bounded to the motion warping of SMPL skeleton in the same way as the human skin is. Rooted in the aforementioned scenarios, SMPL fails to accurately fit such 3D human models.
@@
To this end, we break through the limitations of modeling complicated 3D human gaussians by maintaining an extended joint tree from SMPL skeleton. Although existing SMPL model has the potential capability to manually customize additional skeleton information, time-consuming and case-by-case adjustments are necessary in this case. To overcome this issue, we extend the SMPL skeleton with additional joints for each individual case adaptively. The growth is performed in an explicit and adaptive manner, and enables the fitting of complicated 3D human avatars, offering high-quality rendering with hand-held objects and loose-fitting clothing (e.g., novel view synthesis results in Fig. 1).
@@
The main challenge of extending the SMPL skeleton is to determine where and how to grow additional joints. To avoid the unnecessary memory consumption and potential overfitting caused by an arbitrary growth, we first determine which of the joints are supposed to serve as parent joints by a localization strategy. We have empirically observed that parent joints requiring growth witness larger backpropagation gradients in their associated gaussians due to underfitting. However, determining the association of gaussians with different joints is non-trivial, as the SMPL’s LBS blending weights may not work with the human with hand-held objects or clothing. To more precisely define such association, we introduce the concept of Motion Kernels based on rigid body priors and combine them with LBS weights, resulting in more accurate gradient-based localization. After growth, we adaptively maintain an extended joint tree and update the extra joints by optimizing two MLP decoders for joints positions and rotations. The proposed method, termed as ToMiE, allows for explicit rendering and animation of complicated human avatars represented by extended joints.
@@
By experiments on complicated cases of the DNA-Rendering dataset [7], ToMiE exhibits state-of-the-art rendering quality while maintaining the animatability that is significant for downstream productions. To summarize, our contributions are three-fold as follows:
@@
    ToMiE, a method for creating an enhanced SMPL joint tree via an adaptive growth strategy, which is able to decouple complicated parts from the human body, thereby achieving state-of-the-art results in both rendering and explicit animation on target cases;
@@ 2.1 SMPL-based Human Mesh Avatars
Most of the recent success in digital human modeling can be attributed to the contributions of the SMPL [36] series, which parameterizes the human body as individual shape components and motion-related human poses through 3D mesh scanning and PCA. The pose blend shapes in SMPL describe human body deformations as a linear weighted blending of different joint poses, significantly improving the efficiency of editing and animating digital humans. Furthermore, it has been widely adopted for human body animation, thanks to the methods [9, 49] of estimating SMPL parameters from 2D inputs. Despite their wide range of applications, SMPL and its family still suffer from inherent limitations. Since the originally scanned 3D meshes are skin-tight, according to which the pose blend shapes are learned, the model is unable to handle significantly outlying meshes, such as human with hand-held objects and clothing like skirts. Although this could potentially be solved by auto-rigging [20, 42, 55, 1, 28, 54], it requires 2D/3D shapes as inputs, which is not accessible in human videos.
@@ 2.2 Neural Representation for 3D Human
Methods based on neural representations, such as NeRF [38] and 3DGS [25], have also been playing an important part in digital human reconstruction for their high-quality rendering capabilities. Early NeRF-based methods [52, 45, 27, 6, 13, 3, 4, 5, 10, 11, 12] aim to reconstruct human avatars by inputting monocular or multi-view synchronized videos. [50] enforce smooth priors based on neural Signed Distance Function (SDF) to obtain more accurate human geometry. Recent breakthroughs [32, 35, 60, 30, 17, 46, 29, 18, 40, 24, 16, 31, 34, 57, 26, 39, 21, 58] rely on 3DGS, enabling faster and more accurate rendering. All these methods register the T-pose in a canonical space and use LBS weights to guide the rigid transformations.
2.3 Rendering & Editing of Intricate 3D Human
We revisit the methods attempting to implicitly improve modeling of complicated 3D human. Animatable NeRF [44] defines a per-frame latent code to capture appearance variations across each frame. Simulation methods [48, 2] physically construct simple clothing but struggle with complex one and objects. [29, 14] additionally register global latent bones to compensate for the limitations in clothing rendering. They fail to explicitly decouple the clothing from human body, making precise control infeasible. Another stream [6, 18] leverages a human poses sequence as contexts to resolve appearance ambiguities. But correspondingly, they require a sequence of human poses for animation, adding to the challenges of editing. Moreover, these methods struggle to fit the object-level pose independent of the human poses sequence, such as hand-held items. We also note that some works [15, 37] introduce diffusion-based generative methods to enhance the realism of garment rendering, but these methods are restricted by the traditional SMPL and overlook the editing of complex clothing. Although SMPLicit [8] enables implicit interpolation of clothing types, it remains dependent on SMPL’s LBS process to generate the observed mesh. This limitation prevents localized explicit animation of loose-fitting clothing and restricts its application to external objects. Rendering-based human reconstruction method [19, 33, 53, 23, 22, 57, 59] can not achieve animation. Our approach enables explicit decoupling of hand-held objects and clothing with the human body by extending the SMPL joint tree, allowing for high-quality rendering and explicit animation in complicated scenarios.
@@ 3.1 SMPL(-X) Revisited
The main issue with this LBS-based prior model is that it can only model tight-fitting avatars conforming to the SMPL(-X) paradigm, making it unsuitable for modeling complex human clothing in more generic cases. What even more challenging is its inability to handle hand-held objects that are fully decoupled from the human pose.
@@ 3.2 Human Gaussians Revisited
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (3)
@@ 4 Methods
Figure 2: The pipeline of ToMiE. We initialize the gaussians in the canonical space with a standard SMPL vertices.  (Sec. 4.4) We apply Linear Blend Skinning (LBS) to the gaussian position and utilize a network for rotation and scale correction. During the warmup phase, Adaptive LBS only utilizes the original SMPL skeleton. After adaptive growth, it further includes the newly grown external skeleton. Gaussian rasterization and gradients backpropagation.  (Sec. 4.1, Sec. 4.2) We employ a gradient-based parent joints localization method, and a motion kernel to optimize the gradient assignment process.  (Sec. 4.3) We maintain an extra joint book with MLPs, which generates explicit human pose, enabling the decoupling and explicit animating of hand-held objects and loose-fitting clothing.
Fig. 2 illustrates ToMiE’s adaptive joint growth and gaussian training strategy. Our goal is to extend the SMPL skeleton to handle complicated human scenarios. However, the abuse of growth can lead to unnecessary computational and memory overheads. For an efficient adaptive growth, we first propose a localization strategy of parent joints to ensure that only necessary joints are grown. Furthermore, we explicitly define the hand-held joints in the $SE(3)$ space and optimize them end-to-end through an MLP, ensuring alignment with the original skeleton of SMPL. The extended skeleton can thus support rendering and explicit animation. To address the limitation of LBS in guiding gaussian attributes apart from the positions, we further fine-tune the rotation and scale during training with a deformation field to achieve better non-rigid warping. Next, we will elaborate on these modules and the training strategy.
@@ 4.1 Motion Kernels-guided Joint Gradient Accumulation
A quantitative metric needs to be identified to determine whether a joint requires growth. We notice that, due to the poor fitting ability of existing SMPL-based human gaussian, it will leave larger gradients in complex human regions (e.g., the hand-held object in Fig. 2 ). In other words, identifying joints with larger accumulated gradients can help to indicate which parent joints are more likely to require growth. Bounded with each gaussian, the accumulated gradients will be first assigned to their corresponding joints. Let $g=\|(g_{x},g_{y},g_{z})\|_{2}$ represent the L2 norm of the gradient at each gaussian position. The gradients accumulation $g_{J}$ for the $k$-th joint can then be computed according to
@@
It is worth emphasizing that $\omega_{k}$ here is a weight term determining the assignment of a gaussian to the $k$-th joint. Under the paradigm of SMPL representation, this weight corresponds to the LBS weight $\omega_{\text{lbs0}}$, and guides the rigid transformation of vertices on the SMPL mesh according to the human pose. To account for the differences between the human mesh with clothing and the vanilla SMPL mesh, [17] calculate $\omega_{\text{lbs}}$ by keeping using the LBS weight prior $\omega_{\text{lbs0}}$ and adding an extra learnable network $\Phi_{\text{lbs}}$ for fine-tuning. This formulation (with index $k$ omitted) can be summarized as
@@
where $\text{NN}_{1}$ stands for top-1 nearest-neighbor search algorithm and $\bm{V}$ represents the canonical standard SMPL vertices. It will be, however, clarified by us, that this nearest neighbor-based assigning method will no longer be feasible in cases with handheld objects and complex clothing.
@@ 4.4 Inference Process and Training Strategy
In this subsection, we explain how our adaptive growth is integrated with the training process of human gaussians.
@@
Once the warm-up iterations completes, we begin the adaptive joint growth. With the MK calculated during the warm-up phase, we can locate the parent joints $\bm{J}_{s}$ that require growth. Then we add grown joints to the extra joint book $B^{e}$, optimizing its positions $\bm{j}^{e}$ and rotations $\bm{r}^{e}$ during the subsequent training. Notably, $\Phi_{\text{lbs}}$ needs to extend its output dimensions to include the blending weights for extra joints, following $K=K^{0}+K^{e}$. Since the extra joints do not have prior LBS weights, their blending weights are entirely learned through $\Phi_{\text{lbs}}$.
@@ 5.1 Dataset
Figure 4: Qualitative comparison on the DNA-Rendering dataset [7] with animatable baselines. We show two cases of hand-held objects (0800_07) and loose-fitting clothing (0811_06) (from top to bottom). Im4d* [33] achieves high-quality rendering but cannot be animated, and we compare it in the supplemental materials. Please check the supplemental video for better visualization.
Our method focuses on hand-held objects and loose-fitting clothing, so we select datasets for experiments accordingly. We notice that the DNA-Rendering dataset [7], by capturing complex scenes of the human body, meets our requirements. Specifically, we selected 8 cases that align with our hypothesis, namely 0041_10, 0090_06, 0176_07, 0800_07 (hand-held objects) and 0007_04, 0014 06, 0051_09, 0811_06 (loose-fitting clothing). For each case, we use 24 surrounding views for training and 6 novel surrounding views for testing. All views are synchronized and contained 100 frames each.
In addition to tackle complicated scenarios, it is essential to ensure the model’s performance in typical scenarios involving tight-fitting clothes. Therefore, we additionally test our method on the ZJU-MoCap [45] dataset. Although the tight-clothing cases are too simple to require joint extension, our overall framework still achieves optimal results. Since this part of the experiment is not directly related to adaptive growth, we refer the readers to the supplemental material for further visualizations.